# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v1

@@ Abstract
A common task arising in various domains is that of ranking items based on the outcomes of pairwise comparisons, from ranking players and teams in sports to ranking products or brands in marketing studies and recommendation systems. Statistical inference-based methods such as the Bradley Terry model, which extract rankings based on an underlying generative model of the comparison outcomes, have emerged as flexible and powerful tools to tackle the task of ranking in empirical data. In situations with limited and/or noisy comparisons, it is often challenging to confidently distinguish performance of different items based on the evidence available in the data. However, existing inference-based ranking methods overwhelmingly choose to assign each item to a unique rank or score, suggesting a meaningful distinction when there is none. Here we address this problem by developing a principled Bayesian methodology for learning partial rankings—rankings with ties—that distinguishes among the ranks of different items only when there is sufficient evidence available in the data. Our framework is adaptable to any statistical ranking method in which the outcomes of pairwise observations depend on the ranks or scores of the items being compared. We develop a fast agglomerative algorithm to perform Maximum A Posteriori (MAP) inference of partial rankings under our framework and examine the performance of our method on a variety of real and synthetic network datasets, finding that it frequently gives a more parsimonious summary of the data than traditional ranking, particularly when observations are sparse.
@@ I Introduction
Given the prominence of rankings across domains, it has long been of interest to develop algorithms that automatically rank items based on some set of data about each item whelan2017prior ; davidson1973bayesian . An extensively studied subset of ranking problems is that of ranking from pairwise comparisons langville2012s ; david1963method ; bradley1952rank ; jerdee2024luck , where rankings are inferred from comparisons among pairs of entities. In this approach, the outcome of a single interaction between two items—for example, a win or loss in sports or one product being preferred over another—serves as the basis for the ranking. This approach allows one to infer rankings in situations where underlying ratings of the entities are unavailable, unobserved, or difficult to compute. For example, instead of trying to assess the absolute skill of individual tennis players, we can infer their relative rankings from their head-to-head match results, assuming that matches are more often won by the better player.
@@
In the classic Bradley Terry model bradley1952rank , the outcome of a comparison between two nodes is modeled as a Bernoulli random variable with a probability that depends on the latent scores associated with the items. After inference with the model, one obtains a set of continuous scores that can be ordered to obtain a final ranking. The Bradley Terry model has inspired numerous extensions agresti2012categorical ; jerdee2024luck to more flexibly model variations in match outcomes, including the possibility of ties in the match outcomes rao1967ties ; davidson1970extending ; yan2016ranking ; baker2021modifying (rather than ties in the final rankings as we study here). Another popular inference-based method for ranking from pairwise comparisons is SpringRank de2018physical , which ranks nodes by finding the ground state of a physical system which models the comparison network as a set of directed springs and has also inspired numerous generalizations iacovissi2022interplay ; della2024model . These methods can parsimoniously model heterogeneity and noise in match outcomes but do not allow for the possibility of ties in the inferred rankings, and, as such, are unable to determine when there is enough structure in the comparison data to justify distinguishing the rankings of different items. (In principle, two players can have identical scores under such methods, but this is rarely ever achieved in practice numerically.) Such methods can therefore infer rankings that overfit the data when limited observations are available, resulting in rankings that are highly sensitive to small changes in outcomes. In the ordered Stochastic Block Model (OSBM) of Peixoto peixoto2022ordered , nodes are grouped according to both the mixing structure of the network and hierarchies among the nodes as evidenced by edge directionality. Although the OSBM does allow nodes to be grouped in part by the edge hierarchy, it is heavily influenced by the mixing structure of the nodes, which largely depends on the positions of the edges rather than their directionality (the latter aspect being of primary interest for pairwise ranking). In contrast, the method we propose here focuses only on the directionality of the observed edges and so is adaptable to a broad class of ranking methods, including most popular generalizations of the Bradley Terry model jerdee2024luck .
@@
In this work we introduce a statistical inference-based ranking framework that intrinsically incorporates partial rankings—rankings where multiple nodes can have the same rank—into its underlying generative model. This is accomplished by introducing a hierarchy of uniform Bayesian priors for the underlying rankings of the nodes, which encourages parsimony in the final scores or ranks inferred using the model. This allows us to account for the uncertainty introduced by sparsity in the data by grouping players into the same rank when the data does not provide enough evidence to separate them. Furthermore, our method can be extended to accommodate any inference-based ranking method that models comparison outcomes as a function of the underlying scores or ranks of the nodes involved  bradley1952rank ; de2018physical ; jerdee2024luck . We provide a fast nonparametric agglomerative algorithm to infer the partial rankings according to our method, and by fitting our model to a wide range of synthetic and real-world datasets, we find that it often provides a more parsimonious description of the data in the regime of sparse observations. As a case study, we apply our method to a faculty hiring network among Computer Science departments at US universities clauset2015systematic . The picture that emerges is that of a well-separated hierarchy dominated by a small group of elite universities whose rankings are not meaningfully different statistically, with very little upward mobility across the ranking groups.
@@
The paper is organised as follows. In Sec. II, we introduce the Bradley-Terry model, our partial rankings model, and the agglomerative algorithm we use to perform MAP inference with our model. In Sec. III.1 and Sec. III.2, we apply our method to a wide range of synthetically generated datasets as well as a corpus of empirical networks, finding that partial rankings can provide a more parsimonious description of the data in cases where the networks of outcomes are not sufficiently dense. In Sec. III.3, we focus on a network of faculty hiring among computer science departments in the U.S. and show that our algorithm can be used to extract meaningful ambiguities in the rankings of these departments. We finalize in Sec. IV with our conclusions.
@@ II Methods
Figure 1: Partial rankings in a small example network. (a) Rankings inferred by the Bradley-Terry model and (b) the partial rankings algorithm, for a dataset capturing hierarchical relationships in a pack of wolves. The nodes are labeled according to their inferred ranking. The distance along the y-axis and the node colors are proportional to the inferred strength $\pi_{i}$ of each node $i$, with the strongest nodes placed at the top. In this case, there was not enough statistical evidence in the edges to justify separating the node ranks $\{2,3\}$, $\{6,7,8,9\}$, and $\{11,12\}$ on the left hand side, so the partial ranking method grouped these nodes together into the same partial rankings.
@@ II.1 Bayesian Ranking and the Bradley-Terry Model
Allowing the gradient with respect to $\bm{\pi}$ to vanish gives a self-consistent set of equations for the Maximum Likelihood estimates of the scores. Then, these scores are converted to an implied ranking $\bm{r}$ by simply taking the ordering of the nodes with respect to the inferred strengths $\bm{\pi}$ zermelo1929berechnung ; bradley1952rank ; ford1957solution ; newman2023efficient .
Note that Eq. 2, and therefore Eq. 3, is invariant under a constant rescaling of all player strengths. To resolve this ambiguity, one generally imposes a suitable normalisation. A common choice is to fix the average player score to zero
@@
This choice centres the log-strengths around zero, making them interpretable as relative to an average player. In turn, this ensures that the strength of the average player is given by $\langle\pi\rangle=1$, so that a player with $\pi_{i}>1$ is stronger than average, and one with $\pi_{i}<1$ is weaker than average. This normalisation choice also has the advantage of providing an intuitive interpretation of the player’s strengths. Following newman2023efficient , let $p_{1}$ be the probability that a player with strength $\pi$ has of beating the average player of strength one. Then, according to the Bradley-Terry model, $p_{1}$ is given by $p_{1}=\pi/(\pi+1)$, which in turn implies that $\pi=p_{1}/(1-p_{1})$. A player’s strength is then simply the odds that the player has of beating the average player.
Maximum Likelihood Estimation (MLE) is not the only method for inferring player strengths. Often a prior $P(\bm{\pi})$ is placed on the strengths $\bm{\pi}$ so that a full posterior distribution over $\bm{\pi}$ can be inferred whelan2017prior ; newman2023efficient ; davidson1973bayesian ; caron2012efficient . A common choice for this prior is to assume a uniform distribution over the probability $p_{1}$ that a player with strength $\pi$ defeats the average player, $P(p_{1})=1$. By probability density transformation rules, this is equivalent to setting a logistic prior on the scores $s_{i}=\log\pi_{i}$. Indeed,
@@
$P(\bm{\pi})=\prod_{i}\frac{\pi_{i}}{(\pi_{i}+1)^{2}}.$ (6)
@@
Allowing the gradient with respect to $\bm{\pi}$ to vanish gives a self-consistent set of equations for the MAP estimates $\bm{\hat{\pi}}$, which can be solved in the same way as for the MLEs. The inferred strengths are then converted into an implied ranking $\bm{r}$ by ordering the nodes with respect to the inferred strengths $\bm{\hat{\pi}}$.
@@ II.2 Partial Rankings
As discussed in Sec. I, a limitation of standard MAP estimation for the BT model is that it lacks a mechanism to handle partial rankings. In principle, one could apply a heuristic to assign the same rank to nodes $i$ and $j$ when their scores satisfy $|s_{i}-s_{j}|<\epsilon$ for some small $\epsilon$. However, this approach does not offer any statistical justification for grouping $i$ and $j$ together (i.e., whether the observed score difference is simply due to statistical noise). Alternatively, one could apply 1D numerical clustering to the final score values, but this approach would not account for the significance of score differences in the context of the model’s likelihood or prior.
A principled approach to handling partial rankings in the Bradley-Terry model is to extend the Bayesian hierarchy for the prior on the strengths $P(\bm{\pi})$ by making the strength of each node a function of its underlying ranking $\bm{r}$. This results in the following new MAP objective:
@@
Given Eq. II.2, we now need to define the functional forms of the likelihood and priors. The easiest choice is to use the Bradley-Terry likelihood of Eq. 3, although any choice of likelihood is possible to incorporate into this MAP estimation objective so long as it is a function of latent strengths—i.e., $P(\bm{W}|\bm{\pi})$—or ranks—i.e., $P(\bm{W}|\bm{r})$ (the latter allowing us to remove the prior $P(\bm{\pi}|\bm{r})$). This flexibility allows for the inclusion of a wide range of ranking objectives including extensions of Bradley Terry models jerdee2024luck and SpringRank de2018physical . For our experiments in Sec. III, we will proceed with using the Bradley Terry likelihood of Eq. 3 for subsequent analysis, which facilitates an efficient optimization scheme and a simple comparison with the original Bradley Terry model.
For the prior on $\bm{r}$, we choose to be agnostic with respect to both the number and sizes of the underlying rankings, This results in a hierarchy of uninformative priors, as described by the following generative model:
@@
    Draw the number of unique ranks $R$ uniformly at random from the range $[1,N]$.
@@
    Draw the histogram $\bm{n}=[n_{1},...,n_{R}]$ of the number of nodes of each rank uniformly from the set of all histograms compatible with the number of ranks $R$.
@@
    Draw the ranks $\bm{r}$ uniformly over the set of rankings compatible with the rank counts $\bm{n}$.
@@
where $\Theta$ is the Heaviside step function and $P(\sigma_{r})$ is given by
@@
as in the Bradley-Terry model.
The MAP estimation objective can now be written as the following minimisation problem in terms of the log posterior
@@
denoting the number of edges going from nodes of rank $r$ to nodes of rank $r^{\prime}$ (i.e. the number of times nodes of rank $r$ beat nodes of rank $r^{\prime}$). By minimizing Eq. II.2, we can determine the optimal ranks $\bm{\hat{r}}$ and the corresponding optimal strengths $\bm{\hat{\sigma}}=\{\hat{\sigma}_{\hat{r}_{i}}\}_{i=1}^{N}$ that best fit the observed data.
We can see that the first term in Eq. II.2 is a constant independent of the underlying node ranking and the second and third terms in Eq. II.2 will tend to penalise having a large number of rankings as they will increase $\mathcal{L}(\bm{r},\bm{\sigma})$. The final two terms are harder to interpret, as they will depend on the particular realisation of the matches. However, empirically, we observe that the fourth term will also penalise having a large number of rankings, while the last term will tend to favour them.
An advantage of working with a Bayesian framework is that it provides us with a principled way to compare results obtained by different models. Suppose we have two different ranking models, which we call $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, that correspond to two different hypotheses for the generating mechanism that produced some observed network of outcomes $\bm{W}$. Let $\bm{r}_{1}$ and $\bm{r}_{2}$ be the corresponding most likely rankings obtained by maximising their respective posterior distributions. A principled approach to decide which of the two rankings/model combinations better represents the data is to compute the posterior odds ratio jaynes2003probability
@@
where a ratio above (below) $1$ indicates that we should favour model $\mathcal{H}_{1}$ ($\mathcal{H}_{2}$) based on posterior probability. If we assume that the two models are a priori equally likely ahead of observing any data so that $P(\mathcal{H}_{1})=P(\mathcal{H}_{2})$, we have that the posterior odds ratio can be written as

### v2

@@ Abstract
A common task arising in various domains is that of ranking items based on the outcomes of pairwise comparisons, from ranking players and teams in sports to ranking products or brands in marketing studies and recommendation systems. Statistical inference-based methods such as the Bradley-Terry model, which extract rankings based on an underlying generative model of the comparison outcomes, have emerged as flexible and powerful tools to tackle the task of ranking in empirical data. In situations with limited and/or noisy comparisons, it is often challenging to confidently distinguish the performance of different items based on the evidence available in the data. However, existing inference-based ranking methods overwhelmingly choose to assign each item to a unique rank or score, suggesting a meaningful distinction when there is none. Here, we address this problem by developing a principled Bayesian methodology for learning partial rankings—rankings with ties—that distinguishes among the ranks of different items only when there is sufficient evidence available in the data. Our framework is adaptable to any statistical ranking method in which the outcomes of pairwise observations depend on the ranks or scores of the items being compared. We develop a fast agglomerative algorithm to perform Maximum A Posteriori (MAP) inference of partial rankings under our framework and examine the performance of our method on a variety of real and synthetic network datasets, finding that it frequently gives a more parsimonious summary of the data than traditional ranking, particularly when observations are sparse.
@@ I Introduction
Given the prominence of rankings across domains, it has long been of interest to develop algorithms that automatically rank items based on some set of data about each item bradley1952rank ; mallows1957non ; luce1959individual ; plackett1975analysis ; langville2012s . An extensively studied subset of ranking problems is that of ranking from pairwise comparisons whelan2017prior ; davidson1973bayesian ; langville2012s ; david1963method ; bradley1952rank ; jerdee2024luck , where rankings are inferred from comparisons among pairs of entities. In this approach, the outcome of a single interaction between two items—for example, a win or loss in sports or one product being preferred over another—serves as the basis for the ranking. This approach allows one to infer rankings in situations where underlying ratings of the entities are unavailable, unobserved, or difficult to compute. For example, instead of trying to assess the absolute skill of individual tennis players, we can infer their relative rankings from their head-to-head match results, assuming that matches are more often won by the better player.
@@
In the classic Bradley-Terry (BT) model bradley1952rank , the outcome of a comparison between two nodes is modeled as a Bernoulli random variable with a probability that depends on the latent scores associated with the items. After inference with the model, one obtains a set of continuous scores that can be ordered to obtain a final ranking. The BT model has inspired numerous extensions agresti2012categorical ; jerdee2024luck to more flexibly model variations in match outcomes, including the possibility of ties in the match outcomes rao1967ties ; davidson1970extending ; yan2016ranking ; baker2021modifying (rather than ties in the final rankings as we study here). Another popular inference-based method for ranking from pairwise comparisons is SpringRank de2018physical , which ranks nodes by finding the ground state of a physical system which models the comparison network as a set of directed springs and has also inspired numerous generalizations iacovissi2022interplay ; della2024model . These methods can parsimoniously model heterogeneity and noise in match outcomes but do not allow for the possibility of ties in the inferred rankings and, as such, are unable to determine when there is enough structure in the comparison data to justify distinguishing the rankings of different items. (In principle, two players can have identical scores under such methods, but this is rarely ever achieved in practice numerically.) Such methods can therefore infer rankings that overfit the data when limited observations are available, resulting in rankings that are highly sensitive to small changes in outcomes. Several methods have been proposed to reduce this overfitting and promote ties in the inferred rankings (see Sec. II.2). However, these models require extensive parameter tuning, a priori knowledge about the number and size of the groups being inferred, or incorporate information other than the outcomes of the pairwise comparisons, making them difficult to adapt to a broad range of existing ranking methods.
@@
In this work, we introduce a statistical inference-based ranking framework that intrinsically incorporates partial rankings—rankings where multiple nodes can have the same rank—into its underlying generative model.  [^1] In the context of ranking algorithms, the term partial ranking can generally refer either to rankings where only a subset of the items are fully ordered pearce2024bayesian ; vanichbuncha2017modelling , or where some of the items are tied in ranking fagin2004comparing . In this work, we explicitly refer to partial rankings as rankings with ties. This is accomplished by introducing a hierarchy of uniform Bayesian priors for the underlying rankings of the nodes, which encourages parsimony in the final scores or ranks inferred using the model. This allows us to account for the uncertainty introduced by sparsity in the data by grouping players into the same rank when the data does not provide enough evidence to separate them. Furthermore, our method can be extended to accommodate any inference-based ranking method that models comparison outcomes as a function of the underlying scores or ranks of the nodes involved  bradley1952rank ; de2018physical ; jerdee2024luck . We provide a fast nonparametric agglomerative algorithm to infer the partial rankings according to our method, and by fitting our model to a wide range of synthetic and real-world datasets, we find that it often provides a more parsimonious description of the data in the regime of sparse observations. As a case study, we apply our method to a faculty hiring network among Computer Science departments at U.S. universities clauset2015systematic . The picture that emerges is that of a well-separated hierarchy dominated by a small group of elite universities whose rankings are not meaningfully different statistically, with very little upward mobility across the ranking groups.
@@
The paper is organized as follows. In Sec. II.1 and Sec. II.2, we introduce the BT model and review literature relating to partial ranking inference. In Sec. II.3 and Sec. II.4 and we introduce our partial rankings model and the agglomerative algorithm we use to perform MAP inference with our model. In Sec. III.1 and Sec. III.2, we apply our method to a wide range of synthetically generated datasets as well as a corpus of empirical networks, finding that partial rankings can provide a more parsimonious description of the data in cases where the networks of outcomes are not sufficiently dense. In Sec. III.3, we focus on a network of faculty hiring among computer science departments in the U.S. and show that our algorithm can be used to extract meaningful ambiguities in the rankings of these departments. We finalize in Sec. IV with our conclusions.
@@ II.1 Bayesian Ranking and the Bradley-Terry Model
Differentiating the log-likelihood (i.e. the logarithm of Eq. 3) with respect to the $\pi_{i}$ and equating the result to zero, we obtain the following set of self-consistent equations:
$\pi_{i}=\frac{\sum_{j=1}^{N}w_{ij}}{\sum_{j=1}^{N}(w_{ij}+w_{ji})/(\pi_{i}+\pi_{j})},$ (4)
which can be solved by iteration to obtain the MLE estimates of the player scores. Then, these scores can be converted to an implied ranking $\bm{r}$ by simply taking the ordering of the nodes with respect to the inferred strengths $\bm{\pi}$ zermelo1929berechnung ; bradley1952rank ; ford1957solution ; newman2023efficient . This iterative procedure is known as Zermelo’s algorithm zermelo1929berechnung . More generally, Zermelo’s algorithm has been shown to belong to a larger family of iterative algorithms of the form
$\pi_{i}=\frac{\sum_{j=1}^{N}w_{ij}(\alpha\pi_{i}+\pi_{j})/(\pi_{i}+\pi_{j})}{\sum_{j=1}^{N}(\alpha w_{ij}+w_{ji})/(\pi_{i}+\pi_{j})},$ (5)
which have all been proven to converge to the same MLE solution (if one exists) for $0\leq\alpha\leq 1$. Furthermore, it has been shown that the convergence of this family of algorithms becomes monotonically slower with increasing $\alpha$, with $\alpha=0$ converging fastest and Zermelo’s algorithm (corresponding to the $\alpha=1$ case) being the slowest to converge. Eq. 5 was first derived in newman2023efficient , and we use its $\alpha=0$ realization throughout this work.
The MLE approach described above presents a series of drawbacks. First, we note that Eq. 2, and therefore Eq. 3, is invariant under a constant rescaling of all player strengths, meaning that player scores are not uniquely identifiable. To resolve this ambiguity, one generally imposes a suitable normalization. A common choice is to fix the average player score to zero
@@
This choice centers the log-strengths around zero, making them interpretable as relative to an average player. In turn, this ensures that the strength of the average player is given by $\langle\pi\rangle=1$, so that a player with $\pi_{i}>1$ is stronger than average, and one with $\pi_{i}<1$ is weaker than average. This normalisation choice also has the advantage of providing an intuitive interpretation of the player’s strengths. Following newman2023efficient , let $p_{1}$ be the probability that a player with strength $\pi$ has of beating the average player of strength one. Then, according to the BT model, $p_{1}$ is given by $p_{1}=\pi/(\pi+1)$, which in turn implies that $\pi=p_{1}/(1-p_{1})$. A player’s strength is then simply the odds that the player has of beating the average player.
More problematic is the fact that Eq. 5—and all MLE methods for that matter—have been shown to converge only under the stringent condition that the underlying network of matches is strongly connected, i.e. there exists a directed path between any pair of nodes in the network zermelo1929berechnung ; ford1957solution ; hunter2004mm . If this is not the case, then no maximum of the likelihood function exists, and the scores in Eq. 5 diverge.
Both of these issues can be addressed by adopting a Bayesian perspective and imposing a prior $P(\bm{\pi})$ on the strengths $\bm{\pi}$ so that a full posterior distribution over $\bm{\pi}$ can be inferred whelan2017prior ; newman2023efficient ; davidson1973bayesian ; caron2012efficient . A common choice for this prior is to assume a uniform distribution over the probability $p_{1}$ that a player with strength $\pi$ defeats the average player, $P(p_{1})=1$. By probability density transformation rules, this is equivalent to setting a logistic prior on the scores $s_{i}=\log\pi_{i}$. Indeed,
@@
$P(\bm{\pi})=\prod_{i=1}^{N}\frac{\pi_{i}}{(\pi_{i}+1)^{2}}.$ (8)
The resulting posterior probability for the scores is then given (up to a multiplicative constant) by
$P(\bm{\pi}|\bm{W})\propto\prod_{i,j=1}^{N}\left(\frac{\pi_{i}}{\pi_{i}+\pi_{j}}\right)^{w_{ij}}\prod_{i=1}^{N}\frac{\pi_{i}}{(\pi_{i}+1)^{2}}.$ (9)
Eq. 9 removes the invariance under a rescaling of the scores. Furthermore, it also prevents the scores from diverging. As pointed out in whelan2017prior , the prior for an individual $i$ can be written as
$\frac{\pi_{i}}{(\pi_{i}+1)^{2}}=\frac{\pi_{i}}{\pi_{i}+1}\cdot\frac{1}{\pi_{i}+1},$ (10)
which corresponds to the probability that player $i$ plays two matches against the average player, winning one and losing the other. Eq. 9 can then be seen as the likelihood of a BT model where two pseudo-games have been added for each player, one won and one lost. The resulting network of interactions is now strongly connected, ensuring that the maximum of the posterior always exists, irrespective of whether the original interaction network was fully connected or not.
@@
Again, differentiating with respect to the $\pi_{i}$ and equating the result to zero, we obtain the following set of self-consistent equations for the MAP estimates of the strengths:
$\hat{\pi}_{i}=\frac{1+\sum_{j\neq i}w_{ij}\pi_{j}/(\pi_{i}+\pi_{j})}{2/(\pi_{i}+1)+\sum_{j\neq i}w_{ji}/(\pi_{i}+\pi_{j})},$ (12)
which is the generalization of Eq. 5 (with $\alpha=0$) to the MAP setting and can be solved in the same way as for the MLEs. The primary advantage of Eq. 12 over the MLE updates is that the iterative updates described by Eq. 12 are guaranteed to converge, and do so quite efficiently in practice newman2023efficient . The inferred strengths are then converted into an implied ranking $\bm{r}$ by ordering the nodes with respect to the inferred strengths $\bm{\hat{\pi}}$.
@@ II.2 Partial Rankings
As discussed in Sec. I, a limitation of standard MAP estimation for the BT model is that it lacks a mechanism to handle partial rankings. In principle, one could apply a heuristic to assign the same rank to nodes $i$ and $j$ when their scores satisfy $|s_{i}-s_{j}|<\epsilon$ for some small $\epsilon$. However, this approach does not offer any statistical justification for grouping $i$ and $j$ together (i.e., whether the observed score difference is simply due to statistical noise). Alternatively, one could apply 1D numerical clustering to the final score values, but this approach would not account for the significance of score differences in the context of the model’s likelihood or prior. More principled approaches have been developed for 1D numerical clustering of the scores by applying $L1$ regularization to the adjacent score differences masarotto2012ranking , analogous to the fused lasso tibshirani2005sparsity . This ranking lasso has been extended in a number of different ways to increase its flexibility for different applications fan2001variable ; zou2006adaptive ; masarotto2012ranking ; tutz2015extended ; vana2016computing ; varin2016statistical ; tutz2015extended ; jeon2018sparse ; tutz2015extended ; hermes2024joint .
Despite their widespread application, regularized ranking models suffer from significant drawbacks. The most obvious is the need to tune the lasso penalty parameter, which requires us to fix the level of sparsity a-priori, forcing the groupings in a predefined way and complicating the interpretation of the inferred parameters. Furthermore, as mentioned previously, maximum likelihood estimates for the Bradley-Terry model (and general Bradley-Terry-Luce family of models) without a score prior can be difficult and computationally expensive to perform, although considerable work has been dedicated to developing efficient algorithms hunter2004mm ; maystre2015fast ; agarwal2018accelerated ; nguyen2023efficient ; newman2023efficient . Also, as pointed out in pearce2024bayesian , uncertainty quantification can be hard to perform in lasso-based methods fan2001variable , and prior knowledge on the number or size of the rank groups cannot be incorporated.
Many of these issues can be resolved by resorting to Bayesian approaches. However, these approaches generally require Monte Carlo estimation, which can be slow for estimation davidson1973bayesian ; caron2012efficient . In pearce2024bayesian , Pearce and Erosheva propose a Bayesian method based on the use of a spike-and-slab prior mitchell1988bayesian ; george1997approaches ; ishwaran2005spike to induce parameter fusion. Unlike other methodologies, Pearce and Erosheva’s model requires neither the specification of the number or sizes of the rank cluster nor the tuning of lasso-type penalties, while exploiting conjugacy relationships for estimation of the full posterior. However, this model still requires the selection and tuning of a series of hyperparameters, which influence how aggressively the model fuses the rankings. Furthermore, the particular choice of priors, while facilitating computation and providing strong regularization, might imbue unwarranted structure into the model, particularly in the sparse case, where the data might be insufficient to allow the prior to wash away.
An alternative approach is to formulate fully nonparametric models, which do not require any hyperparameter tuning. In the ordered Stochastic Block Model (OSBM) peixoto2022ordered , Peixoto proposes a nonparametric model based on a modification of the Stochastic Block Model holland1983stochastic ; peixoto2019bayesian in which nodes are grouped according to both the mixing structure of the network and hierarchies among the nodes as evidenced by edge directionality. Although the OSBM does allow nodes to be grouped in part by the edge hierarchy, it is heavily influenced by the mixing structure of the nodes, which largely depends on the positions of the edges rather than their directionality (the latter aspect being of primary interest for pairwise ranking). In contrast, the method we propose here focuses only on the directionality of the observed edges and so is adaptable to a broad class of ranking methods, including most popular generalizations of the BT model jerdee2024luck .
II.3 Bayesian Nonparametric Model
A principled approach to handling partial rankings is to extend the Bayesian hierarchy for the prior on the strengths $P(\bm{\pi})$ by making the strength of each node a function of its underlying ranking $\bm{r}$. This results in the following new MAP objective:
@@ II.3 Bayesian Nonparametric Model
Given Eq. II.3, we now need to define the functional forms of the likelihood and priors. The easiest choice is to use the BT likelihood of Eq. 3, although any choice of likelihood is possible to incorporate into this MAP estimation objective so long as it is a function of latent strengths—i.e., $P(\bm{W}|\bm{\pi})$—or ranks—i.e., $P(\bm{W}|\bm{r})$ (the latter allowing us to remove the prior $P(\bm{\pi}|\bm{r})$). This flexibility allows for the inclusion of a wide range of ranking objectives including extensions of BT models jerdee2024luck and SpringRank de2018physical . For our experiments in Sec. III, we will proceed with using the BT likelihood of Eq. 3 for subsequent analysis, which facilitates an efficient optimization scheme and a simple comparison with the original BT model.
For the prior on $\bm{r}$, we choose to be agnostic with respect to both the number and sizes of the underlying rankings. This results in a hierarchy of uniform priors, as described by the following generative model:
@@
    Draw the number of unique ranks $R$ uniformly at random from the range $[1,N]$. As there are $N$ maximum possible unique ranks, the probability of observing any value of $R\in[1,N]$ is given by
    $P(R)=\frac{1}{N}.$ (14)
@@
    Draw the histogram $\bm{n}=[n_{1},...,n_{R}]$ of the number of nodes of each rank uniformly from the set of all histograms compatible with the number of ranks $R$. This amounts to selecting a list of $R$ positive integer values that sum to $N$. As there are ($N-1\choose R-1$ such lists, each draw occurs with probability
    $P(\bm{n}|R)=\frac{1}{{N-1\choose R-1}}.$ (15)
@@
    Draw the ranks $\bm{r}$ uniformly over the set of rankings compatible with the rank counts $\bm{n}$. Since there are ($N\choose n_{1},...,n_{R}$ possible partitions of $N$ objects into $R$ groups such that each group $r$ has size $n_{r}$, the total probability of this step is given by
    $P(\bm{r}|\bm{n})=\frac{1}{{N\choose n_{1},...,n_{R}}}.$ (16)
@@
where $\Theta$ is the Heaviside step function (which ensures that the unique strengths are ordered) and $P(\sigma_{r})$ is given by
@@
as in the BT model.
The MAP estimation objective can now be written as the following minimization problem in terms of the (negative) log posterior
@@
denoting the number of edges going from nodes of rank $r$ to nodes of rank $r^{\prime}$ (i.e. the number of times nodes of rank $r$ beat nodes of rank $r^{\prime}$). By minimizing Eq. II.3, we can determine the optimal ranks $\bm{\hat{r}}$ and the corresponding optimal strengths $\bm{\hat{\sigma}}=\{\hat{\sigma}_{\hat{r}_{i}}\}_{i=1}^{N}$ that best fit the observed data. It is important to emphasize that, although this work focuses primarily on MAP estimates of the strengths and rankings, Eq. II.3 defines a well-posed (negative) log-posterior distribution. This posterior can be explored using sampling techniques such as Markov Chain Monte Carlo (MCMC), enabling estimation of the full posterior distribution and associated uncertainty quantification rather than relying solely on point estimates.
We can see that the first term in Eq. II.3 is a constant independent of the underlying node ranking. The second and third terms synergistically penalize having a large number of rankings as they will increase $\mathcal{L}(\bm{r},\bm{\sigma})$. Indeed, it is easy to prove that the sum of these two terms
$f(R)=\log\binom{N-1}{R-1}+\log\binom{N}{n_{1}\ldots n_{R}}$ (24)
has a global minimum when $R=1$, see Appendix A. The final two terms are harder to interpret, as they will depend on the particular realization of the matches. However, empirically, we observe that the fourth term will also penalize having a large number of rankings, while the last term will tend to favor them.
An advantage of working with a Bayesian framework is that it provides us with a principled way to compare results obtained by different models. Suppose we have two different ranking models, which we call $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, that correspond to two different hypotheses for the generating mechanism that produced some observed network of outcomes $\bm{W}$. Let $\bm{r}_{1}$ and $\bm{r}_{2}$ be the corresponding most likely rankings obtained by maximizing their respective posterior distributions. A principled approach to decide which of the two rankings/model combinations better represents the data is to compute the posterior odds ratio jaynes2003probability
@@
where a ratio above (below) $1$ indicates that we should favor model $\mathcal{H}_{1}$ ($\mathcal{H}_{2}$) based on posterior probability. If we assume that the two models are a priori equally likely ahead of observing any data so that $P(\mathcal{H}_{1})=P(\mathcal{H}_{2})$, we have that the posterior odds ratio can be written as